What is four consecutive integers whose sum is 194
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



"What are four consectutive integers..." (Yes, correct grammar counts)

Let , , , and represent the 4 integers. Then:



Solve for , then calculate , , and .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

four consecutive integers whose sum is 194
-----
1st: x-2
2nd: x-1
3rd: x
4th: x+1
----
Equation:
sum = 194
4x -2 = 194
4x = 196
x = 49
---
1st: x-2 = 47
2nd: x-1 = 48
3rd: x = 49
4th: x+1 = 50
=======================
Cheers,
Stan H.
================